I am pushing a shopping cart, causing it to accelerate at a rate of 0.18 m/s2.
I would like to double my rate of acceleration. How much harder must I push?

I am pushing a shopping cart, causing it to accelerate at a rate of 0.16 m/s2.
If this shopping cart had more groceries in it, causing it to be twice as massive, how much more would it accelerate?

A 9.8 kg object is moving at a constant speed of 1.5 m/s.
What is this object's acceleration? What is the net force acting on this object?

I pull a 2.5 kilogram block of wood across a table with a force of 19.6 Newtons.
It slides at a constant speed.
What is the weight of this block of wood? What is the normal force acting on this block of wood?  What is the acceleration of this block of wood?  What is the net force on this block of wood?   What is the strength (or magnitude) of the frictional force on this block of wood?

If I pull a 1.7 kilogram block of wood across the table with a force of 15 Newtons, and the force of friction between the block and the table is 5 Newtons, what is the net force acting on this block?

What is the acceleration of this block of wood?

An airplane's engines provide 15 tons of thrust to keep it moving at a steady velocity.
How much air drag is acting on the airplane?

A 2890 kilogram car is cruising down the straight, level road at a constant speed.
What is this car's weight?
What is the normal force acting on this car?
What is the net force acting on this car?
What is this car's acceleration?

If the driver wants to accelerate at a rate of 2 m/s2, how much additional force must the car apply against the road?

Two people are pushing a car. Jack pushes with a force of 270 Newtons. Klement helps by pushing in the same direction with a force of 160 Newtons.
What is the net force on this car?
Tommy comes along and pushes in the opposite direction with a force of 96 Newtons. What is the net force acting on the car now?

A 53 kilogram girl is falling to Earth using a parachute.
She has reached terminal velocity.
What is this girl's weight?
What is the net force acting on this girl?
What is the force of the air drag on the girl?
Answer #1

Solving the first 4 questions

1)

Since Force = mass x acceleration, hence for acceleration to double either the mass should be halved or the force should be made double

Therefore if F1/m = a = 0.18

and 2a = 0.36 = F2/m

hence F2 = 2 * F1 hence twice harder must be pushed.

2)

Assuming the force to be constant we get m1 * a1 = m2 * a2

m2 = 2 * m1

hence a1 = 2 * a2

hence a2 = 0.16 / 2 = 0.08 m/s^2

3)

Since the object is moving with constant velocity hence the acceleration of the body = 0 thus force = 0.

4)